# Modular Control of Biological Networks

### v1

@@ Abstract
The concept of control is central to understanding and applications of biological network models. Some of their key structural features relate to control functions, through gene regulation, signaling, or metabolic mechanisms, and computational models need to encode these. Applications of models often focus on model-based control, such as in biomedicine or metabolic engineering. This paper presents an approach to model-based control that exploits two common features of biological networks, namely their modular structure and canalizing features of their regulatory mechanisms. The paper focuses on intracellular regulatory networks, represented by Boolean network models. A main result of this paper is that control strategies can be identified by focusing on one module at a time. This paper also presents a criterion based on canalizing features of the regulatory rules to identify modules that do not contribute to network control and can be excluded. For even moderately sized networks, finding global control inputs is computationally very challenging. The modular approach presented here leads to a highly efficient approach to solving this problem. This approach is applied to a published Boolean network model of blood cancer large granular lymphocyte (T-LGL) leukemia to identify a minimal control set that achieves a desired control objective.
${}^{1}$epartment of Mathematics, University of Kentucky, Lexington, KY 40506, USA
${}^{2}$epartment of Mathematics, University of Dayton, Dayton, Ohio 45469, USA
${}^{3}$athematics Department, California Polytechnic State University, San Luis Obispo, CA 93407, USA
${}^{4}$epartment of Mathematics, Iowa State University, Ames, IA 50011, USA
${}^{5}$epartment of Medicine, University of Florida, Gainesville, FL 32610, USA
@@ 1 Introduction
For the commonly-used modeling framework of ordinary differential equations, there is a well-developed theory of optimal control, which is largely absent for other modeling frameworks, such as Boolean networks or agent-based models, both frequently used in systems biology and biomedicine. Furthermore, control inputs, in many cases, are of a binary nature, such as gene knockouts or the blocking of mechanisms. For BNs, there is no readily available mathematical theory that could be used for control, leaving sampling and simulation. As networks get larger, with hundreds [7] or even thousands of nodes [8], this leaves few computational tools to identify control inputs for achieving preselected objectives, such as moving a network from one phenotype (e.g., cancer) to another (e.g., normal). One approach is to reduce the system in a way that the reduced system maintains relevant dynamical properties such as its attractors [9, 10]. This allows the control methods to be applied to the reduced system, and the same controls can then be used for the original system.
@@
An approach that has not been used so far, to our knowledge, is to exploit the modular structure of many biological systems to identify control strategies by focusing on one module at a time. Modularity refers to the division of the system into separate units, or modules, that each have a specific function [11, 12]. Modularity is a fundamental property of biological systems that is essential for the evolution of new functions and the development of robustness [13, 14]. In [15], we developed a mathematical theory of modularity for Boolean network models and showed that one can identify network-level control inputs at the modular level. That is, we obtain global control inputs by identifying them at the local, modular level and assembling them to global control. This enables network control for much larger networks than would otherwise be computationally feasible. In this paper, we develop this approach into a mathematical theory of biological network control.
@@
We further propose to use another property of biological networks, represented through Boolean network models. Almost all Boolean rules that describe the dynamics of over 120 published, expert-curated biological Boolean network models have the property that they exhibit some degree of canalization [16]. A Boolean function is canalizing if it has one or more variables that, when they take on a particular value, they determine the value of the function, irrespective of the values of all the other variables. As an example, any variable in a conjunctive rule (e.g., $x\cap y\cap z$) determines the value of the entire rule, when it takes on the value 0. We derive a criterion for Boolean network models whose Boolean functions are all canalizing, that can be used to exclude certain modules from needing to be considered for the identification of controls.
@@
Figure 1: Control via modularity. First, the network is decomposed into its constituent modules: $F_{1},\dots,F_{n}$. Then, controls $\mu_{1},\dots,\mu_{n}$ are identified for each module. Combining the controls of the modules $\mu=(\mu_{1},\ldots,\mu_{n})$ yields a set of controls for the whole network.
@@ 2.1 Boolean Networks
The wiring diagram of a Boolean network $F=(f_{1},\ldots,f_{n}):\operatorname{\mathbb{F}}_{2}^{n}\rightarrow\operatorname{\mathbb{F}}_{2}^{n}$ is the directed graph with vertices $x_{1},\ldots,x_{n}$ and an edge from $x_{i}$ to $x_{j}$ if $f_{j}$ depends on $x_{i}$. That is, if there exists $\mathbf{x}\in\operatorname{\mathbb{F}}_{2}^{n}$ such that $f_{j}(\mathbf{x})\neq f(\bf x+e_{i}),$ where $e_{i}$ is the $i$th unit vector.
@@
Figure 2: Wiring diagram and state space of the Boolean network in Example 2.1-2.8. (a) The wiring diagram encodes the dependency between variables. (b) The state space is a directed graph with edges between all states and their images. This graph therefore encodes all possible trajectories.
@@ 2.2 Dynamics of Boolean networks
Another directed graph associated with a BN is the state space. It describes all possible transition of the BN from one time step to another. The attractors of a BN are sets of states from which there is no escape as the system evolves. An attractor with a single state is also called a steady state (or fixed point). In mathematical models of intracellular regulatory networks, the attractors of the model are often associated with the possible phenotypes of the cell. This idea can be traced back to Waddington [18] and Kauffman [19]. For example, in a model of cancer cells, the steady states of the model correspond to proliferative, apoptotic, or growth-arrest phenotypes [20]. Mathematically, a phenotype is associated with a group of attractors where a subset of the system’s variables have the same states. These shared states are then used as biomarkers that indicate diverse hallmarks of the system.
@@
The space of attractors for a Boolean network is the collection $\mathcal{D}(F)$ of all minimal subsets $\mathcal{C}\subseteq\operatorname{\mathbb{F}}_{2}^{n}$ satisfying $F(\mathcal{C})=\mathcal{C}$.
@@
    The subset $\mathcal{D}^{1}(F)\subset\mathcal{D}(F)$ of sets of exact size 1 consists of all steady states (also known as fixed points) of $F$.
@@
    The subset $\mathcal{D}^{r}(F)\subset\mathcal{D}(F)$ of sets of exact size $r$ consists of all cycles of length $r$ of $F$.
@@ 2.3 Modules
Before defining modules we will describe the restriction of a Boolean network to a subset of its variables using the following example. For a formal definition of restrictions, please see [15].
@@
with wiring diagram in Figure 3a. The restriction of this network to $Y=\{x_{1},x_{2}\}$ is the 2-variable network $F|_{Y}(x_{1},x_{2})=(x_{2}\wedge x_{1},\neg x_{1})$, which forms the first module (indicated by the amber box in Figure 3a), while the restriction of $F$ to $Y=\{x_{3},x_{4}\}$ is the 2-variable network $F|_{Y}(x_{3},x_{4})=(\neg x_{4},x_{3}\wedge x_{4})$, which forms the second module (indicated by the green module in Figure 3a). Note that the wiring diagram of $F|_{Y}$ is always a subgraph of the wiring diagram of $F$, irrespective of the choice of $Y$.
Figure 3: Boolean network decomposition into modules. (a) Wiring diagram of a non-strongly connected Boolean network where the non-trivial modules are highlighted by amber and green boxes. (b) Directed acyclic graph describing the corresponding connections between the nontrivial modules.
Definition 2.10.
Remark 2.11.
The wiring diagram of any Boolean network is either strongly connected or it is composed of a collection of strongly connected components where connections between different component move in only one direction.
Let $F$ be a Boolean network and let $W_{1},\ldots,W_{m}$ be the strongly connected components of its wiring diagram, with $X_{i}$ denoting the set of variables in strongly connected component $W_{i}$. Then, the modules of $F$ are defined by $F|_{X_{1}},\ldots,F|_{X_{n}}$.
Definition 2.12.
Let $W_{1},\ldots,W_{m}$ be the strongly connected components of the wiring diagram of a Boolean network $F$. By setting $W_{i}\rightarrow W_{j}$ if there exists at least one edge from a vertex in $W_{i}$ to a vertex in $W_{j}$, we obtain a (directed) acyclic graph
$Q=\{(i,j)|W_{i}\rightarrow W_{j}\},$
which describes the connections between the strongly connected components of $F$.
Example 2.13.
For the Boolean network $F$ from Example 2.9, the wiring diagram has two strongly connected components $W_{1}$ and $W_{2}$ with variables $X_{1}=\{x_{1},x_{2}\}$ and $X_{2}=\{x_{3},x_{4}\}$ (Figure 3a), connected according to the directed acyclic graph $Q=\{(1,2)\}$. The two modules of $F$ are given by the restriction of $F$ to $X_{1}$ and $X_{2}$, that is, $F|_{X_{1}}(x_{1},x_{2})=(x_{2}\wedge x_{1},\neg x_{1})$ and $F|_{X_{2}}(x_{3},x_{4})=(\neg x_{4},x_{3}\wedge x_{4})$ (Figure 3b). Note that the module $F|_{X_{1}}$, i.e., the restriction of $F$ to $X_{1}$, is simply the projection of $F$ onto the variables $X_{1}$ because $W_{1}$ does not receive feedback from the other component (i.e., because $(2,1)\not\in Q$).
@@ 3 Control via Modularity
Once the modules have been identified, different methods for phenotype control (that is, control of the attractor space) can be used to identify controls in these networks. Some of these methods employ stable motifs [23], feedback vertex sets [24], as well as algebraic approaches [25, 26, 27]. For our examples below, we will use the methods defined in [23, 25, 24] to find controls for the simple networks.
A Boolean network $F=(f_{1},\ldots,f_{n}):\operatorname{\mathbb{F}}_{2}^{n}\rightarrow\operatorname{\mathbb{F}}_{2}^{n}$ with control is a Boolean network $\mathcal{F}:\mathbb{F}_{2}^{n}\times U\rightarrow\mathbb{F}_{2}^{n}$, where $U$ is a set that denotes all possible controls, defined below. The case of no control coincides with the original Boolean network, that is, $\mathcal{F}(x,0)=F(x)$. Given a control $u\in U$, the dynamics are given by $x(t+1)=\mathcal{F}(x(t),u)$. See [25] for additional details and examples of how to encode control edges and nodes in a Boolean network.
@@
$\mathcal{F}_{j}(x,u_{i,j}):=f_{j}(x_{1},\dots,(u_{i,j}+1)x_{i}+u_{i,j}a_{i},\dots,x_{n}),$ (1)
where $a_{i}$ is a constant in $\operatorname{\mathbb{F}}_{2}$, encodes the control of the edge $x_{i}\rightarrow x_{j}$, since for each possible value of $u_{i,j}\in\mathbb{F}_{2}$ we have the following control settings:
  • If $u_{i,j}=0$, $\mathcal{F}_{j}(x,0)=f_{j}(x_{1},\dots,x_{i},\dots,x_{n})$. That is, the control is not active.
  • If $u_{i,j}=1$, $\mathcal{F}_{j}(x,1)=f_{j}(x_{1},\dots,x_{i}=a_{i},\dots,x_{n})$. In this case, the control is active, and the action represents the removal of the edge $x_{i}\rightarrow x_{j}$ when $a_{i}=0$, and the constant expression of the edge if $a_{i}=1$. We use $x_{i}\xrightarrow[]{a_{i}}x_{j}$ to denote that the control is active.
This definition can be easily extended for the control of many edges, so that we obtain $\mathcal{F}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{e}\rightarrow\mathbb{F}_{2}^{n}$, where $e$ is the number of edges in the wiring diagram. Each coordinate, $u_{i,j}$, of $u$ in $\mathcal{F}(x,u)$ encodes the control of an edge $x_{i}\rightarrow x_{j}$.
@@
$\mathcal{F}_{j}(x,u^{-}_{i},u^{+}_{i}):=(u^{-}_{i}+u^{+}_{i}+1)f_{j}(x)+u^{+}_{i}$ (2)
encodes the control (knock-out or constant expression) of the node $x_{i}$, since for each possible value of $(u^{-}_{i},u^{+}_{i})\in\mathbb{F}_{2}^{2}$ we have the following control settings:
  • For $u^{-}_{i}=0,u^{+}_{i}=0$, $\mathcal{F}_{j}(x,0,0)=f_{j}(x)$. That is, the control is not active.
  • For $u^{-}_{i}=1,u^{+}_{i}=0$, $\mathcal{F}_{j}(x,1,0)=0$. This action represents the knock-out of the node $x_{j}$.
  • For $u^{-}_{i}=0,u^{+}_{i}=1$, $\mathcal{F}_{j}(x,0,1)=1$. This action represents the constant expression of the node $x_{j}$.
  • For $u^{-}_{i}=1,u^{+}_{i}=1$, $\mathcal{F}_{j}(x,1,1)=f_{j}(x_{t_{1}},\dots,x_{t_{m}})+1$. This action changes the Boolean function to its negative value.
@@
A control $\mu$ stabilizes a network $F$ at an attractor $\mathcal{C}$ when the resulting network after applying $\mu$ to $F$ (denoted here as $F^{\mu}$) has $\mathcal{C}$ as its only attractor.
For a Boolean network $F$, we let $\mathcal{D}(F)$ denote the set of its attractors. Whenever the Boolean network $F$ is decomposable into multiple constituent modules $F_{1}$, $F_{2}$, $\cdots$, $F_{n}$ ($n\geq 2$), we write $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{n-1}}F_{n}$ where the semi-product operation $\rtimes_{P_{i}}$ indicates the coupling of the subnetworks, as described in [15]. Furthermore, from the decomposition theory described in [15], the attractors of $F$ are of the form $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\cdots\oplus\mathcal{C}_{n}$ where $\mathcal{C}_{i}\in\mathcal{D}(F_{i})$ is an attractor of the subnetwork, for $i=1,\dots,n$. The following theorem takes advantage of the modular structure of the network to find controls one module at a time.
@@
Given a decomposable network $F=F_{1}\rtimes_{P}F_{2}$, if $\mu_{1}$ is a control that stabilizes $F_{1}$ in $\mathcal{C}_{1}$ (whether $\mathcal{C}_{1}$ is an existing attractor or a new one) and $\mu_{2}$ is a control that stabilizes $F^{\mathcal{C}_{1}}_{2}$ in $\mathcal{C}_{2}$ (whether $\mathcal{C}_{2}$ is an existing attractor or a new one), then $\mu=(\mu_{1},\mu_{2})$ is a control that stabilizes $F$ in $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$ provided that either $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state.
@@
Let $F^{\mu_{1}}_{1}$ be the resulting network after applying the control $\mu_{1}$. Thus, the dynamics of $F^{\mu_{1}}_{1}$ is $\mathcal{C}_{1}$, that is $\mathcal{D}(F^{\mu_{1}}_{1})=\mathcal{C}_{1}$. Similarly, the dynamics of $F^{\mathcal{C}_{1},\mu_{2}}_{2}$ is $\mathcal{C}_{2}$. That is, $\mathcal{D}(F^{\mathcal{C}_{1},\mu_{2}}_{2})=\mathcal{C}_{2}$. Then,

### v3

# Modular control of Boolean network models

@@ Abstract
The concept of control is crucial for effectively understanding and applying biological network models. Key structural features relate to control functions through gene regulation, signaling, or metabolic mechanisms, and computational models need to encode these. Applications often focus on model-based control, such as in biomedicine or metabolic engineering. In a recent paper, the authors developed a theoretical framework of modularity in Boolean networks, which lead to a canonical semidirect product decomposition of these systems. In this paper, we present an approach to model-based control that exploits this modular structure, as well as the canalizing features of the regulatory mechanisms. We show how to identify control strategies from the individual modules, and we present a criterion based on canalizing features of the regulatory rules to identify modules that do not contribute to network control and can be excluded. For even moderately sized networks, finding global control inputs is computationally challenging. Our modular approach leads to an efficient approach to solving this problem. We apply it to a published Boolean network model of blood cancer large granular lymphocyte (T-LGL) leukemia to identify a minimal control set that achieves a desired control objective.
1Department of Mathematics, University of Kentucky, Lexington, KY 40506, USA
2Department of Mathematics, University of Dayton, Dayton, OH 45469, USA
3Mathematics Department, California Polytechnic State University, San Luis Obispo, CA 93407, USA
4Department of Mathematics, Iowa State University, Ames, IA 50011, USA
5Department of Medicine, University of Florida, Gainesville, FL 32610, USA
@@ 1 Introduction
For the commonly-used modeling framework of ordinary differential equations, there is a well-developed theory of optimal control, which is largely absent from other modeling frameworks, such as Boolean networks or agent-based models, both frequently used in systems biology and biomedicine. Furthermore, control inputs, in many cases, are of a binary nature, such as gene knockouts or the blocking of mechanisms. Existing Boolean network control methods do not scale well. As networks get larger, with hundreds [7] or even thousands of nodes [8], only few computational tools can identify control inputs for achieving preselected objectives, such as moving a network from one phenotype (e.g., cancer) to another (e.g., normal). One approach is to reduce the system in a way that the reduced system maintains relevant dynamical properties such as its attractors [9, 10]. This allows the control methods to be applied to the reduced system, and the same controls can then be used for the original system.
Control targets in Boolean networks have been identified by a variety of approaches: using stable motifs [11, 12], feedback vertex sets [13, 14], trap spaces [15, 16], model checking [17], and other methods [18, 19]. A few further approaches have explicitly used strongly connected components (SCCs) for model analysis by decomposing the wiring diagram or the state space of the network [20, 21]. For example, stable motifs describe a set of network nodes and their corresponding states which are such that the nodes form a minimal SCC (e.g. a feedback loop) and their states form a partial fixed point of the Boolean network [22, 11]. An identification of the stable motifs yields not only an efficient way to derive the network attractors but also control targets. Trap spaces, defined in [23] and related to stable motifs, are subspaces of the entire state space of a Boolean network that the dynamics cannot escape. Trap spaces, which can often be efficiently computed for biological Boolean models, have recently been used to simplify the identification of both network attractors and controls [15, 16]. However, to our knowledge, none of the approaches developed thus far exploit the modular structure exhibited by many biological systems, in order to identify control strategies by focusing on one module at a time, which is the approach used in this paper. This approach can substantially simplify the problem of identifying suitable controls, specifically in networks that consist of multiple decently-sized modules. For example, we showed in [24] that our modular approach can efficiently solve the control identification problem for a 69-node pancreatic cancer model that consists of three modules of size greater than one, plus a number of single-node modules. Some of the existing control methods can often not handle reasonably large models of e.g. 70 nodes, as has been shown in [4, 25]. It is important to note, however, that our control approach does not guarantee finding a control of minimal size or all possible controls that achieve a given objective.
@@
Modularity refers to the division of the system into separate units, or modules, that each have a specific function [26, 27]. Modularity is a fundamental property of biological systems that is essential for the evolution of new functions and the development of robustness [28, 29]. In [24], we developed a mathematical theory of modularity for Boolean network models and showed that one can identify network-level control inputs at the modular level. That is, we obtain global control inputs by identifying them at the local, modular level and assembling them to global control. This enables network control for much larger networks than would otherwise be computationally unfeasible. It is worth noting that, although the ideas and concepts behind our approach to modularity are intuitive and natural, to our knowledge, we developed the first rigorous mathematical theory of modular decomposition of Boolean networks, which we expand to questions related to the identification of biological network controls in this paper.
@@
We further propose to use another property of biological networks, represented through Boolean network models. Almost all Boolean rules that describe the dynamics of over 120 published, expert-curated biological Boolean network models have the property that they exhibit some degree of canalization [30]. A Boolean function is canalizing if it has one or more variables that, when they take on a particular value, they determine the value of the function, irrespective of the values of all the other variables. As an example, any variable in a conjunctive rule (e.g., $x\wedge y\wedge z$) determines the value of the entire rule, when it takes on the value 0. We derive a criterion for Boolean network models whose Boolean functions are all canalizing, that can be used to exclude certain modules from needing to be considered for the identification of controls.
@@
Figure 1: Control via modularity. First, the network is decomposed into its constituent modules: $F_{1},\dots,F_{m}$. Then, controls $\mu_{1},\dots,\mu_{m}$ are identified for each module. Combining the controls of the modules $\mu=(\mu_{1},\ldots,\mu_{m})$ yields a set of controls for the whole network.
@@ 2.1 Boolean Networks
Definition 2.1.
The wiring diagram of a Boolean network $F=(f_{1},\ldots,f_{n}):\operatorname{\mathbb{F}}_{2}^{n}\rightarrow\operatorname{\mathbb{F}}_{2}^{n}$ is the directed graph with vertices $x_{1},\ldots,x_{n}$ and an edge from $x_{i}$ to $x_{j}$ if $f_{j}$ depends on $x_{i}$. That is, if there exists $\mathbf{x}\in\operatorname{\mathbb{F}}_{2}^{n}$ such that $f_{j}(\mathbf{x})\neq f_{j}(\bf x+e_{i}),$ where $\bf e_{i}$ is the $i$th unit vector.
Definition 2.2.
Remark 2.3.
The wiring diagram of any Boolean network is either strongly connected or it is composed of a collection of strongly connected components where connections between different components move in only one direction.
@@
Figure 2: Wiring diagram and state space of the Boolean network in Example 2.4-2.11. (a) The wiring diagram encodes the dependency between variables. (b) The state space is a directed graph with edges between all states and their images. This graph therefore encodes all possible trajectories.
@@ 2.2 Dynamics of Boolean networks
Another directed graph associated with a BN is the state transition graph, also referred to as the state space. It describes all possible transitions of the BN from one time step to the next. The attractors of a BN are minimal sets of states from which there is no escape as the system evolves. An attractor with a single state is also called a steady state (or fixed point). In mathematical models of intracellular regulatory networks, the attractors of the model are often associated with the possible phenotypes of the cell. This idea can be traced back to Waddington [32] and Kauffman [33]. For example, in a model of cancer cells, the steady states of the model correspond to proliferative, apoptotic, or growth-arrest phenotypes [34]. Mathematically, a phenotype is associated with a group of attractors where a subset of the system’s variables have the same states. These shared states are then used as biomarkers that indicate diverse hallmarks of the system.
@@
The set of attractors for a Boolean network is the set $\mathcal{A}(F)$ of all minimal nonempty subsets $\mathcal{C}\subseteq\operatorname{\mathbb{F}}_{2}^{n}$ satisfying $F(\mathcal{C})=\mathcal{C}$.
@@
    The subset $\mathcal{A}^{1}(F)\subset\mathcal{A}(F)$ of sets of exact size 1 consists of all steady states (also known as fixed points) of $F$.
@@
    The subset $\mathcal{A}^{r}(F)\subset\mathcal{A}(F)$ of sets of exact size $r$ consists of all cycles of length $r$ of $F$.
@@ 2.3 Modules
In [24], a concept of modularity was introduced for Boolean networks. The decomposition into modules occurs on structural (wiring diagram) level but induces an analogous decomposition of the network dynamics, in the sense that one can recover the dynamics of the entire network from the dynamics of the modules. For this decomposition, a module of a BN is defined as a subnetwork that is itself a BN with external parameters in the subset of variables that specifies a strongly connected component (SCC) in the wiring diagram (see Example 2.12). More precisely, for a Boolean network $F$ and subset of its variables $S$, we define the restriction of $F$ to $S$ to be the BN $F|_{S}=(f_{k_{1}},\ldots,f_{k_{m}})$, where $x_{k_{i}}\in S$ for $i=1,\ldots,m$. We note that $f_{k_{i}}$ might contain inputs that are not part of $S$ (e.g., when $x_{k_{i}}$ is regulated by some variables that are not in $S$). Therefore, the BN $F|_{S}$ may contain external parameters (which are themselves fixed and do not possess an update rule). Given a $F$ with wiring diagram $W$, let $W_{1},\ldots,W_{m}$ be the SCCs of $W$ with pairwise disjoint sets of variables $S_{i}$. The modules of $F$ are then the restrictions to these sets of variables, $F|_{S_{i}}$. Further, the modular structure of $F$ can be described by a directed acyclic graph $Q=(V,E)$ with vertex set $V=\{W_{1},\ldots,W_{m}\}$ (where we contracted each SCC into a single vertex) and edge set $E=\{(i,j)\mid W_{i}\longrightarrow W_{j}\}$ by setting $W_{i}\longrightarrow W_{j}$ whenever there exists a node from $W_{i}$ to $W_{j}$ (we show a modular decomposition of a small BN in Example 2.12).
@@
with wiring diagram in Figure 3a. The restriction of this network to $S_{1}=\{x_{1},x_{2}\}$ is the 2-variable network $F|_{S_{1}}(x_{1},x_{2})=(x_{2}\wedge x_{1},\neg x_{1})$, which forms the first module (indicated by the amber box in Figure 3a), while the restriction of $F$ to $S_{2}=\{x_{3},x_{4}\}$ is the 2-variable network with external parameters $x_{1}$ and $x_{2}$ $F|_{S_{2}}(x_{3},x_{4})=(x_{1}\vee\neg x_{4},(x_{1}\wedge\neg x_{2})\vee(x_{3}\wedge x_{4})))$, which forms the second module (indicated by the green module in Figure 3a). Note that the module $F|_{S_{1}}$, i.e., the restriction of $F$ to $S_{1}$, is simply the projection of $F$ onto the variables $S_{1}$ because $W_{1}$ does not receive feedback from the other component.
The wiring diagram of this network has two strongly connected components $W_{1}$ and $W_{2}$ with variables $S_{1}=\{x_{1},x_{2}\}$ and $S_{2}=\{x_{3},x_{4}\}$ (Figure 3a), connected according to the directed acyclic graph $Q=(V,E)$ with vertex set $V=\{W_{1},W_{2}\}$ (where we contracted each SCC into a single vertex) and edge set $E=\{(1,2)\mid W_{1}\longrightarrow W_{2}\}$ (Figure 3b).
Figure 3: Boolean network decomposition into modules. (a) Wiring diagram of a non-strongly connected Boolean network where the modules are highlighted by amber and green boxes. (b) Directed acyclic graph describing the corresponding connections between the modules.
@@ 3 Control via Modularity
Once the modules have been identified, different methods for phenotype control (that is, control of the attractor space) can be used to identify controls in these networks. Some of these methods employ stable motifs [11], feedback vertex sets [13], as well as algebraic approaches [37, 38, 39]. For our examples below, we will use the methods defined in [11, 37, 13] to find controls for the modules.
A Boolean network $F=(f_{1},\ldots,f_{n}):\operatorname{\mathbb{F}}_{2}^{n}\rightarrow\operatorname{\mathbb{F}}_{2}^{n}$ with control is a Boolean network $F^{\mu}:\mathbb{F}_{2}^{n}\times U\rightarrow\mathbb{F}_{2}^{n}$, where $U$ is a set that denotes all possible controls, defined below. The case of no control coincides with the original Boolean network, that is, $F^{\mu}(x,0)=F(x)$. Given a control $\mu\in U$, the dynamics are given by $x(t+1)=F^{\mu}(x(t),\mu)$. See [37] for additional details and examples of how to encode control edges and nodes in a Boolean network.
@@
$F^{\mu}_{j}(x,\mu_{i,j}):=f_{j}(x_{1},\dots,(\mu_{i,j}+1)x_{i}+\mu_{i,j}a_{i},\dots,x_{n}),$ (1)
where $a_{i}$ is a constant in $\operatorname{\mathbb{F}}_{2}$, encodes the control of the edge $x_{i}\rightarrow x_{j}$, since for each possible value of $\mu_{i,j}\in\mathbb{F}_{2}$ we have the following control settings:
  • If $\mu_{i,j}=0$, $F^{\mu}_{j}(x,0)=f_{j}(x_{1},\dots,x_{i},\dots,x_{n})$. That is, the control is not active.
  • If $\mu_{i,j}=1$, $F^{\mu}_{j}(x,1)=f_{j}(x_{1},\dots,x_{i}=a_{i},\dots,x_{n})$. In this case, the control is active, and the action represents the removal of the edge $x_{i}\rightarrow x_{j}$ when $a_{i}=0$, and the constant expression of the edge if $a_{i}=1$. We use $x_{i}\xrightarrow[]{a_{i}}x_{j}$ to denote that the control is active.
This definition can be easily extended for the control of many edges, so that we obtain $F^{\mu}:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{e}\rightarrow\mathbb{F}_{2}^{n}$, where $e$ is the number of edges in the wiring diagram. Each coordinate, $\mu_{i,j}$, of $\mu$ in $F^{\mu}(x,\mu)$ encodes the control of an edge $x_{i}\rightarrow x_{j}$.
@@
$F^{\mu}_{j}(x,\mu^{-}_{i},\mu^{+}_{i}):=(\mu^{-}_{i}+\mu^{+}_{i}+1)f_{j}(x)+\mu^{+}_{i}$ (2)
encodes the control (knock-out or constant expression) of the node $x_{i}$, since for each possible value of $(\mu^{-}_{i},\mu^{+}_{i})\in\mathbb{F}_{2}^{2}$ we have the following control settings:
  • For $\mu^{-}_{i}=0,\mu^{+}_{i}=0$, $F^{\mu}_{j}(x,0,0)=f_{j}(x)$. That is, the control is not active.
  • For $\mu^{-}_{i}=1,\mu^{+}_{i}=0$, $F^{\mu}_{j}(x,1,0)=0$. This action represents the knock-out of the node $x_{j}$.
  • For $\mu^{-}_{i}=0,\mu^{+}_{i}=1$, $F^{\mu}_{j}(x,0,1)=1$. This action represents the constant expression of the node $x_{j}$.
  • For $\mu^{-}_{i}=1,\mu^{+}_{i}=1$, $F^{\mu}_{j}(x,1,1)=f_{j}(x_{t_{1}},\dots,x_{t_{m}})+1$. This action changes the Boolean function to its negative value. This case is usually not considered in the control search since it is biologically impractical to implement.
We note that the algebraic framework is versatile enough that we can encode any type of control, such as a combination of node and edge control at the same time.
@@
A control $\mu$ stabilizes a given Boolean network $F:\operatorname{\mathbb{F}}_{2}^{n}\rightarrow\operatorname{\mathbb{F}}_{2}^{n}$ at an attractor $\mathcal{C}$ $\subseteq\operatorname{\mathbb{F}}_{2}^{n}$ when the resulting network after applying $\mu$ to $F$ (denoted here as $F^{\mu}$) has $\mathcal{C}$ as its only attractor. Note that $\mu$ may contain a combination of one or multiple edge or node controls, as described in Definitions 1 and 3.2.
For a Boolean network $F$, we let $\mathcal{A}(F)$ denote the set of its attractors. Whenever a Boolean network $F$ has more than one module we say that it is decomposable into its constituent modules $F_{1},F_{2},\ldots,F_{m}$ ($m\geq 2$), and write $F=F_{1}\rtimes_{P_{1}}F_{2}\rtimes_{P_{2}}\cdots\rtimes_{P_{m-1}}F_{m}$ where the semi-product operation $\rtimes_{P_{i}}$ indicates the coupling of the subnetworks, as described in [24] (see Example 3.4 for an example of coupling). Furthermore, from the decomposition theory described in [24], the attractors of $F$ are of the form $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\cdots\oplus\mathcal{C}_{n}$ where $\mathcal{C}_{i}\in\mathcal{A}(F_{i})$ is an attractor of the subnetwork, for $i=1,\dots,n$ (one can think of this direct sum operation as concatenating the attractors from the different modules, see Example 3.6 where we show an example of this operation).
Example 3.4.
Consider the Boolean network in Example 2.12,
with wiring diagram in Figure 3a. Let $F_{1}=F|_{S_{1}}(x_{1},x_{2})=(x_{2}\wedge x_{1},\neg x_{1})$ be the restriction of $F$ to $S_{1}=\{x_{1},x_{2}\}$ which forms the first module (indicated by the amber box in Figure 3a) and let $F_{2}=F|_{S_{2}}(x_{3},x_{4})=(e_{1}\vee\neg x_{4},(e_{1}\wedge\neg e_{2})\vee(x_{3}\wedge x_{4})))$ be the restriction of $F$ to $S_{2}=\{x_{3},x_{4}\}$, which forms the second module (indicated by the green box in Figure 3a). Here, $e_{1}$ and $e_{2}$ are external parameters of $F|_{S_{2}}$ that take the place of $x_{1}$ and $x_{2}$, which is indicated by the coupling $P=\{x_{1}\to e_{1},x_{2}\to e_{2}\}$. $F$ is thus decomposable into $F_{1}$ and $F_{2}$ and we have $F=F_{1}\rtimes_{P}F_{2}$.
The following theorem takes advantage of the modular structure of the network to find controls one module at a time.
@@
Given a decomposable network $F=F_{1}\rtimes_{P}F_{2}$, if $\mu_{1}$ is a control that stabilizes $F_{1}$ in $\mathcal{C}_{1}$ (whether $\mathcal{C}_{1}$ is an existing attractor of $F_{1}$ or a new one, after applying control $\mu_{1}$) and $\mu_{2}$ is a control that stabilizes $F_{2}$ in $\mathcal{C}_{2}$ (whether $\mathcal{C}_{2}$ is an existing attractor of $F_{2}$ or a new one, after applying control $\mu_{2}$), then $\mu=(\mu_{1},\mu_{2})$ is a control that stabilizes $F$ in $\mathcal{C}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$ provided that at least one of $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ is a steady state.
@@
Let $F^{\mu_{1}}_{1}$ be the resulting network after applying the control $\mu_{1}$. Thus, the dynamics of $F^{\mu_{1}}_{1}$ is $\mathcal{C}_{1}$, that is $\mathcal{A}(F^{\mu_{1}}_{1})=\mathcal{C}_{1}$. Similarly, the dynamics of $F^{\mu_{2}}_{2}$ is $\mathcal{C}_{2}$. Let $F^{\mathcal{C}_{1},\mu_{2}}_{2}$ be the network that results after setting all external parameters of $F_{2}$ to the states of $\mathcal{C}_{1}$ and applying the control $\mu_{2}$. Then, $\mathcal{A}(F^{\mathcal{C}_{1},\mu_{2}}_{2})=\mathcal{C}_{2}$. Thus,